A plane travels 300 miles against the wind in five hours, then turns and travels 330 miles with the wind in three hours.Find the speed of the wind and the speed of the plane .
Answer by mananth(16946)

Plane speed =x mph
wind speed =y mph
against wind 5 hours
with wind 3 hours

Distance against 300 miles distance with 330 miles
t=d/r against wind -
300.00 / ( x - y )= 5.00
5.00 ( x - y ) = 3.00
5.00 x - 5.00 y = 300.00 ....................1

330.00 / ( x + y )= 3.00
3.00 ( x + y ) = 330.00
3.00 x + 3.00 y = 330.00 ...............2
Multiply (1) by 3.00
Multiply (2) by 5.00
we get
15.00 x + -15.00 y = 900.00
15.00 x + 15.00 y = 1650.00
30.00 x = 2550.00
/ 30.00
x = 85.00 mph

plug value of x in (1) y
5.00 x -5.00 y = 300.00
425.00 -5.00 -425.00 = 300.00
-5.00 y = 300.00
-5.00 y = -125.00 mph
y = 25.00
Plane speed 85.00 mph
wind speed 25.00 mph

m.ananth@hotmail.ca